OpenStudy (anonymous):

HEEEEELP MEDAL if 9^x-9^x-1=24, then find the value of (6x)^x

OpenStudy (anonymous):

first solve (9^x-9^(x-1) = 24) x = 3/2 ((1/2)*(6*3))^(3/2) = ?

OpenStudy (anonymous):

umm 27

OpenStudy (anonymous):

yep!